OpenStudy (anonymous):

If f(x) = −4x − 1, find f(−2).

OpenStudy (hba):

just plugin -2

OpenStudy (hba):

@koala305 please respond

hero (hero):

Hint: x = -2

OpenStudy (anonymous):

thats f(x)=-4*-2-1 =8-1 =7